Consider the reaction 2H2(g) + O2(g) → 2H2O(g).

The following bond energies have been determined experimentally:

  • H-H 436 kJ/mol
  • O=O 497 kJ/mol
  • H-O 464 kJ/mol

Suppose we start with 4 mol of hydrogen and end up with 4 mol of H2O.

If the system was at 300 K before the reaction took place, what is the final temperature after the reaction?
Answer #1

You would find the total bond energy for both sides of the equation, then subtract.

Reactants:

H2: 436 kJ/mol
O2: 498 kJ/mol

Total: 1370 kJ/mol

Products

structured: 2(H-O-H )

2(464 kJ/mol + 464 kJ/mol) = 1856 kJ/mol


The bond energy is greater for the products than the reactants, which actually means that it takes more energy to break the bonds of the products. Therefore, the reaction is exothermic.

1370 kJ/mol - 1856 kJ/mol = -486 kJ/mol

For 4 moles q = 4*486= 1944 KJ

Use the equation for heat transfer to express the heat lost

where is the symbol for heat transfer, is the mass of the substance, and is the change in temperature. The symbol stands for specific heat and depends on the material and phase

1944 = mc(T-300)

Based on the material m and c can be applied and the final temperature can be found out.